# MAPLE: Self-supervised Learning-Enhanced Nonlinear Dimensionality Reduction for Visual Analysis

### v1

@@ III-C3 Step 3: MMCR Loss Computation and Parameter Update
Maximizing the between-variance encourages the centroids of local manifolds to spread out across multiple orthogonal directions. In other words, it promotes variance to be distributed more evenly across multiple principal components, rather than collapsing into a few dominant axes. Geometrically, this pushes centroids to form large pairwise angles on the unit sphere, suppressing concentration in a single direction that would otherwise create hub nodes in the neighborhood graph.
@@
MMCR loss function.  Using the between-variance and within-variance of local manifolds, we define the MMCR loss as:
@@
Efficient optimization.  Although Steps 1–3 could, in theory, be repeated at every update, we adopt a more efficient optimization strategy. First, we use mini-batch gradient descent, computing the loss over batches of data. Second, we reuse the constructed graph across updates to Steps 1 and 3, and only reconstruct it every fixed number of epochs (20 by default).
@@ III-D Phase 2: Graph Layout
where $\tau$ is a softmax temperature parameter and $\sum_{j}w_{ij}=1$. $w_{ij}$ defines a probability distribution over $i$’s neighbors: closer neighbors (higher similarity) receive higher weights, while the softmax normalizes across the neighborhood, balancing the pulling forces from different neighbors. MAPLE’s fuzzy weight computation differs from UMAP’s original one (Eq. 1), which is designed to eliminate weak neighbor connections. Since the learned graph in MAPLE has already removed such connections, applying UMAP’s procedure would result in unnecessary elimination.
@@ III-D2 Step 2: Layout Optimization
Implementation  We implemented MAPLE in Python 3, using NumPy/SciPy [virtanen2020scipy] for matrix calculations and scikit-learn [pedregosa2011scikit] for kkNN graph construction. The neural networks are built with PyTorch [paszke2019pytorch]. For the MMCR loss, we adopted the implementation provided by the original authors [yerxa2023mmcr]. For layout optimization, we employed the official UMAP implementation [mcinnes_umap_2020]. Additional implementation details, including neural network settings, are provided in the supplementary materials.
@@ IV Results
kNN classification accuracy (core) $\uparrow$ 0.813 0.794 0.818 0.778 0.302 0.260 0.650 0.554 0.660 0.555 0.825 0.790 0.859 0.855 0.678 0.601
@@
Calinski–Harabasz $\uparrow$ 9282.101 11480.073 25828.332 36460.773 643.450 522.385 521.646 471.087 557.774 517.456 463.124 457.781 678.806 1205.347 560.435 1262.673
Davies–Bouldin $\downarrow$ 0.926 0.971 2.055 1.947 16.232 34.250 15.664 19.263 8.131 9.393 3.719 3.170 5.360 4.245 5.119 6.050
@@
We evaluate MAPLE on eight datasets with multiple classes or subtypes, representative of real-world challenges. Fig. 3 shows layouts generated by MAPLE and five other DR methods. Tab. I presents quantitative results across 17 evaluation measures on the same datasets (in the same order as Fig. 3), comparing MAPLE with UMAP using a Euclidean input-space metric. The corresponding layouts can be found in the first two rows of Fig. 3.
@@
    PBMC 4K from 10x Genomics [10xgenomics], annotated with the Cellarium Cell Annotation Service [cell_annotation_service]: N=4, [^0] N{=}4{,}000, $D=33,538$
@@ IV-A Quantitative Evaluation
Quantitative evaluation of DR layouts is inherently challenging, as there is no universally accepted “ground truth” for high-dimensional structure, and different measures emphasize different aspects of layout quality [espadoto2019toward, jung2024ghostumap, machado2025necessary, jeon2023classes]. To perform quantitative evaluation, it is therefore necessary to select or design quality measures for DR layouts [jeon2023zadu, jeon2025unveiling]. While existing measures provide valuable insights, selecting appropriate ones remains challenging due to inherent trade-offs in the structural aspects they capture. A large class of existing measures heuristically quantifies how well neighborhood relationships in high-dimensional space are preserved in the resulting layout. For instance, the Kullback–Leibler divergence [hinton2002stochastic] compares neighborhood probability distributions, while trustworthiness and continuity [venna2006local] penalize neighbors that are gained or lost compared to the input kkNN. ZADU [jeon2023zadu], a library compiling various quality measures, includes 19 measures, 15 of which fall into this class.
Although widely adopted as standard evaluation criteria for new DR methods [atzberger2024large, jung2025ghostumap2], these measures capture only a limited aspect of layout quality and face two critical issues. First, many rely on kkNN graphs constructed with distance metrics; however, as discussed in Sec. II-A and Sec. III-A, conventional kkNN graphs do not necessarily reflect intrinsic manifold structure [beyer1999nearest, bergman2020deep, bohm2022unsupervised, wang2025dimension]. Second, these measures emphasize statistical interpretability rather than geometric or informational expressiveness. They quantify whether certain local statistics (e.g., distances among neighbors) are preserved. However, as highlighted in SSL research [bear2020learning, galvez2023role], raw data often contain noise, and such statistics may not reflect whether embeddings and layouts provide faithful geometric representations of the manifold (e.g., curvature, geodesics [moon_PHATE_2019]). MAPLE is designed to address data with noisy local manifolds, and the layout quality should be judged by the ability to reveal semantically meaningful structures.
Given these limitations, we use class labels as a proxy for ground truth semantic structure in high-dimensional space. Accordingly, we use label-based measures to quantify layout quality. Jeon et al. [jeon2023classes] argue that label-based measures can be unreliable as classes do not always correspond to clusters in the original space. This concern is mitigated in our evaluation by our choice of datasets, where classes are well defined.
@@
We design perceptual measures inspired by perceptual similarity tests in computer vision [zhang2018unreasonable], which mimic how people judge whether two images look alike. We adopt the same sampling-based idea for evaluating the layouts perceptually. We randomly sample reference points from the layout produced by MAPLE or UMAP, and compare distances to same-class and different-class neighbors. Based on Zhang et al. [zhang2018unreasonable], we define a two-alternative forced choice (2AFC) accuracy measure, which counts how often a reference is closer to a same-class point than to a different-class point. Similar to the Just Noticeable Difference (JND) test [zhang2018unreasonable], we also compute the area under the receiver operating characteristic curve (ROC-AUC) score [bradley1997use], which corresponds to sampling many pairs of points from the layout, recording their pairwise distances, and labeling them as same-class or different-class. The ROC curve captures how well pairwise distance discriminates between same-class and different-class pairs across thresholds, and its AUC then summarizes and provides a single reliability score. Unlike the original perceptual measure that relies on deep neural network features (e.g., from AlexNet [krizhevsky2012imagenet]) to approximate human visual perception, we sample points on the DR results and check them with ground-truth labels.
@@
T-shirt Trouser Pullover Dress Coat Sandal Shirt Sneaker Bag Boot
@@
Other observations.  Regarding global class separability, MAPLE generally produces layouts with clearer margins between classes compared to UMAP. On global measures, both the Davies–Bouldin [vergani2018soft] and Calinski–Harabasz [calinski1974dendrite] scores compare within-cluster scatter against between-cluster separation, and they favor compact, well-separated clusters. MAPLE exhibits mixed performance compared to UMAP, depending on the data geometry. By revealing fine-grained classes, MAPLE often increases inter-class separation (thereby improving these scores); in some other cases, however, this increased resolution resulted in elongated structures that are penalized by these compactness-based measures. MAPLE also achieves higher Hausdorff distances [huttenlocher2002comparing, taha2015metrics] across all experiments.
Clustering and perceptual measures can assess MAPLE’s quality for downstream analytical workflows such as annotation and cluster analysis. For clustering measures, MAPLE outperforms UMAP on almost all datasets. This confirms that the layouts produced by MAPLE are easier to partition and consistent with the improvement in local measures. For perceptual measures (2AFC, ROC-AUC) [zhang2018unreasonable], MAPLE and UMAP are comparable. The two perceptual measures are ranking-based tests that ask whether same-class distances are consistently smaller than different-class distances, so they measure binary rankings of pairwise distances in the layout. Unless a method achieves substantially better class separation, layouts with rounded and compact clusters that UMAP often produces tend to score higher because their within-class distances are more uniformly small and stable. In contrast, MAPLE produces higher-resolution structures that can introduce greater variance within classes, which can lead to lower perceptual scores for datasets like Fashion-MNIST [xiao2017fashion].
@@ IV-B Qualitative Evaluation
Visual validation.    To examine the patterns revealed by MAPLE and relate them to the underlying images, we designed a grid-based aggregation visualization (see Fig. 4). For an interactively selected region of interest, we discretize the 2D space into a grid. In each grid cell containing data points, we visualize the pixel-wise average of the corresponding images. This analysis indicates that the structures produced by MAPLE correspond to visually meaningful fashion sub-types disentangled by the self-supervised objective. As shown in Fig. 4(a), the upper-left branch in the ‘Trouser’ class primarily contains wide-leg shorts, while the upper-right region captures thin, straight-leg cuts. The central cluster consists of standard jeans, and the bottom peaks correspond to trousers with one bent leg. Similarly, in the region containing the ‘Coat’ and ‘Dress’ classes (Fig. 4(b)), MAPLE shows roughly three distinct subareas: coats at the top, dark clothing to the left, and skirts or dresses (identified by longer lower hems) at the bottom. Additional examples on the MNIST handwritten digits dataset are provided in the supplementary materials.
This ability to resolve intra-class variation is a key strength of MAPLE. By encouraging local neighborhoods to align with lower-rank structures through a self-supervised objective, MAPLE introduces perceptible gaps in the final layout between visually distinct sub-groups. From an information visualization perspective, these gaps align with the Gestalt law of proximity [7294427], supporting the perception of distinct groups and facilitating the identification of subtle data sub-types. While we do not claim that all geometric structures in the layout directly correspond to visually identifiable semantic components, the observed consistency between local layout structures and input-level visual patterns suggests that a substantial portion of the internal structures reflects coherent variation in the data and arises from the objective’s suppression of redundant variance. Overall, these results indicate that MAPLE enables higher-resolution visual exploration and supports the discovery of fine-grained intra-class patterns that are often obscured in other DR methods.
Figure 5: Completion time of MAPLE and UMAP in seconds. Left: runtime with varying feature dimensions ($D$) at fixed $N=10,000$. Right: runtime with varying data sizes (NN) at fixed $D=784$. Solid and dotted lines indicate neighborhood sizes of $k=15$ and $k=30$, respectively. Both axes are shown on logarithmic scales, with tick labels $10^{2}$ and $10^{3}$ indicate orders of magnitude.
@@ IV-C Scalability and Runtime Analysis
We evaluate the performance of MAPLE on a laptop equipped with an Apple M1 Pro chip and 16 GB of unified memory. The training framework is implemented in PyTorch using the Metal Performance Shaders (MPS) backend to execute tensor operations on the integrated GPU. To improve throughput, we use Automatic Mixed Precision (AMP) for gradient computations. Due to current limitations in the MPS backend for certain singular-value decomposition operators, the nuclear norm required by the MMCR loss is computed on the CPU. To accommodate large datasets under limited memory, we use a batch-loading strategy that avoids memory overflow.
@@
Observation.  The results are shown in Fig. 5. MAPLE exhibits approximately linear runtime scaling with respect to both data size $N$ and feature dimensionality $D$, closely matching the scaling behavior of standard UMAP. For instance, increasing the data size from N=50, [^0] N=50{,}000 to N=100, [^0] N=100{,}000 increases runtime by approximately 2.7 times (from roughly 4 minutes to 11 minutes). Runtime also scales linearly with $D$. However, we observe increased sensitivity to neighborhood size for high-dimensional inputs. At D=10, [^0] D=10{,}000, increasing $k$ from 15 to 30 leads to a significant runtime increase (from 261 seconds to 1,925 seconds). This behavior arises from two factors. First, the encoder processes entire local neighborhoods, causing the dominant computational cost to scale as $O(kND^{\prime})$. Second, at high dimensionality, the resulting dense tensors exceed CPU cache capacity, making system memory bandwidth a limiting factor. While batching prevents memory overflow, the runtime bottleneck shifts to data loading and computation.
The runtime difference between MAPLE and UMAP reflects an inherent trade-off in learning-based DR methods. MAPLE incurs additional computational cost due to the neural network training phase, resulting in a runtime of the order of minutes rather than seconds. However, this cost is predictable and scales linearly with dataset size. In practice, MAPLE’s runtime is comparable to other parametric DR methods [sainburg2021parametric]. Moreover, this investment yields a reusable, geometry-aware encoder. Our results show that MAPLE can process datasets with up to 100,000100{,}000 samples on a standard laptop within minutes, without requiring high-performance hardware.

### v2

@@ III-C3 Step 3: MMCR Loss Computation and Parameter Update
Maximizing the between-variance encourages the centroids of local manifolds to spread across multiple orthogonal directions. In other words, it promotes variance to be distributed more evenly across multiple principal components, rather than collapsing into a few dominant axes. Geometrically, this pushes centroids to form large pairwise angles on the unit sphere, suppressing concentration in a single direction that would otherwise create hub nodes in the neighborhood graph.
@@
MMCR loss function.  Using the between- and within-variances of local manifolds, we define the MMCR loss as:
@@
Efficient optimization.  Although Steps 1–3 could, in theory, be repeated at every update, we adopt a computationally efficient optimization strategy. First, we use mini-batch gradient descent, computing the loss over batches of data. Second, we reuse the constructed graph across updates in Steps 1 and 3, and only reconstruct it every fixed number of epochs (20 by default).
@@ III-D Phase 2: Graph Layout
where $\tau$ is a softmax temperature parameter and $\sum_{j}w_{ij}=1$. $w_{ij}$ defines a probability distribution over $i$’s neighbors: closer neighbors (higher similarity) receive higher weights, while the softmax normalizes across the neighborhood, balancing the attractive forces from different neighbors. MAPLE’s fuzzy weight computation differs from UMAP’s original one (Eq. 1), which is designed to eliminate weak neighbor connections. Since the learned graph in MAPLE has already removed such connections, applying UMAP’s procedure would result in unnecessary elimination.
@@ III-D2 Step 2: Layout Optimization
Implementation  We implemented MAPLE in Python 3, using NumPy/SciPy [virtanen2020scipy] for matrix calculations and scikit-learn [pedregosa2011scikit] for kkNN graph construction. The neural networks are built with PyTorch [paszke2019pytorch]. For the MMCR loss, we adopted the implementation by the original authors [yerxa2023mmcr]. For layout optimization, we employed the official UMAP implementation [mcinnes_umap_2020]. Additional implementation details, including neural network settings, are provided in the supplementary materials.
@@ IV Results
kkNN classification accuracy (core) $\uparrow$ 0.813 0.794 0.818 0.778 0.302 0.260 0.650 0.554 0.660 0.555 0.825 0.790 0.859 0.855 0.678 0.601
@@
CH $\uparrow$ 9282.101 11480.073 25828.332 36460.773 643.450 522.385 521.646 471.087 557.774 517.456 463.124 457.781 678.806 1205.347 560.435 1262.673
DB $\downarrow$ 0.926 0.971 2.055 1.947 16.232 34.250 15.664 19.263 8.131 9.393 3.719 3.170 5.360 4.245 5.119 6.050
@@
We evaluate MAPLE on eight datasets with multiple classes or subtypes, representative of real-world high-dimensional analysis challenges. Fig. 3 shows layouts generated by MAPLE and five other DR methods. Tab. I presents quantitative results across 17 evaluation measures on the same datasets (in the same order as Fig. 3), comparing MAPLE with UMAP using a Euclidean input-space metric. The corresponding layouts can be found in the first two rows of Fig. 3.
@@
    PBMC 4K from 10x Genomics [10xgenomics], annotated with the Cellarium Cell Annotation Service [cell_annotation_service]: N=4, [^0] N{=}4{,}000, D=33,538D=33{,}538
@@ IV-A Quantitative Evaluation
Quantitative evaluation of DR layouts is inherently challenging, as there is no universally accepted “ground truth” for high-dimensional structure, and different measures emphasize different aspects of layout quality [espadoto2019toward, jung2025ghostumap2, machado2025necessary, jeon2023classes]. To perform quantitative evaluation, it is therefore necessary to select or design quality measures for DR layouts [jeon2023zadu, jeon2025unveiling]. While existing measures provide valuable insights, selecting appropriate ones remains challenging due to inherent trade-offs in the structural aspects they capture. A large class of existing measures heuristically quantifies how well neighborhood relationships in high-dimensional space are preserved in the resulting layout. For instance, the Kullback–Leibler divergence [hinton2002stochastic] compares neighborhood probability distributions, while trustworthiness and continuity [venna2006local] penalize neighbors that are gained or lost compared to the input kkNNs. ZADU [jeon2023zadu], a library compiling various quality measures, includes 20 measures, 15 of which fall into this class.
Although widely adopted as standard evaluation criteria for new DR methods [atzberger2024large, jung2025ghostumap2], these measures capture only limited aspects of layout quality and face two critical issues. First, many rely on kkNN graphs constructed with distance metrics; however, as discussed in Sec. II-A and Sec. III-A, conventional kkNN graphs do not necessarily reflect intrinsic manifold structure [beyer1999nearest, bergman2020deep, bohm2022unsupervised, wang2025dimension]. Second, these measures emphasize statistical interpretability rather than geometric or informational expressiveness. They quantify whether certain local statistics (e.g., distances among neighbors) are preserved. However, as highlighted in SSL research [bear2020learning, galvez2023role], raw data often contain noise, and such statistics may not reflect whether embeddings and layouts provide faithful geometric representations of the manifold (e.g., curvature, geodesics [moon_PHATE_2019]). MAPLE is designed to address data with noisy local manifolds, and the layout quality should be judged by the ability to reveal semantically meaningful structures.
Given these limitations, we use class labels as a proxy for ground truth semantic structure in high-dimensional space. Accordingly, we use label-based measures to quantify layout quality. Jeon et al. [jeon2023classes] argue that label-based measures can be unreliable as classes do not always correspond to clusters in the original space. This concern is mitigated in our evaluation by our choice of datasets, where classes are well-defined.
@@
We design perceptual measures inspired by perceptual similarity tests in computer vision [zhang2018unreasonable], which mimic how people judge whether two images look alike. We adopt the same sampling-based idea for evaluating the layouts perceptually. We randomly sample reference points from the layout produced by MAPLE or UMAP, and compare distances to same-class and different-class neighbors. Based on Zhang et al. [zhang2018unreasonable], we define a two-alternative forced choice (2AFC) accuracy measure, which counts how often a reference is closer to a same-class point than to a different-class point. Similar to the Just Noticeable Difference (JND) test [zhang2018unreasonable], we also compute the area under the receiver operating characteristic curve (ROC–AUC) score [bradley1997use], which corresponds to sampling many pairs of points from the layout, recording their pairwise distances, and labeling them as same-class or different-class. The ROC curve captures how well pairwise distance discriminates between same-class and different-class pairs across thresholds, and the resulting AUC provides a single reliability score. Unlike the original perceptual measure that relies on deep neural network features (e.g., from AlexNet [krizhevsky2012imagenet]) to approximate human visual perception, we sample points on the DR results and check them with ground-truth labels.
@@
T-shirt Trouser Pullover Dress Coat Sandal Shirt Sneaker Bag Ankle Boot
@@
Other observations.  Regarding global class separability, MAPLE generally produces layouts with clearer margins between classes compared to UMAP. On global measures, both the DB [vergani2018soft] and CH [calinski1974dendrite] scores compare within-cluster scatter against between-cluster separation, and they favor compact, well-separated clusters. MAPLE exhibits mixed performance compared to UMAP, depending on the data geometry. By revealing fine-grained classes, MAPLE often increases inter-class separation (thereby improving these scores); in some other cases, however, this increased resolution resulted in elongated structures that are penalized by these compactness-based measures. MAPLE also achieves higher Hausdorff distances [huttenlocher2002comparing, taha2015metrics] across all experiments.
Clustering and perceptual measures can assess MAPLE’s quality for downstream analytical workflows such as annotation and cluster analysis. For clustering measures, MAPLE outperforms UMAP on almost all datasets. This confirms that the layouts produced by MAPLE are easier to partition and consistent with the improvement in local measures. For perceptual measures (2AFC, ROC–AUC) [zhang2018unreasonable], MAPLE and UMAP are comparable. The two perceptual measures are ranking-based tests that ask whether same-class distances are consistently smaller than different-class distances, so they measure binary rankings of pairwise distances in the layout. Unless a method achieves substantially better class separation, layouts with rounded and compact clusters that UMAP often produces tend to score higher because their within-class distances are more uniformly small and stable. In contrast, MAPLE produces higher-resolution structures that can introduce greater variance within classes, which can lead to lower perceptual scores for datasets like Fashion-MNIST [xiao2017fashion].
@@ IV-B Qualitative Evaluation
Visual validation.    To examine the patterns revealed by MAPLE and relate them to the underlying images, we designed a grid-based aggregation visualization (see Fig. 4). For an interactively selected region of interest, we discretize the 2D space into a grid. In each grid cell containing data points, we visualize the pixel-wise average of the corresponding images. This analysis indicates that the structures produced by MAPLE correspond to visually meaningful fashion subtypes disentangled by the self-supervised objective. As shown in Fig. 4(a), the upper-left branch in the ‘Trouser’ class primarily contains wide-leg shorts, while the upper-right region captures thin, straight-leg cuts. The central cluster consists of standard jeans, and the bottom peaks correspond to trousers with one bent leg. Similarly, in the region containing the ‘Coat’ and ‘Dress’ classes (Fig. 4(b)), MAPLE shows roughly three distinct subareas: coats at the top, dark clothing to the left, and skirts or dresses (identified by longer lower hems) at the bottom. Additional examples on the MNIST handwritten digits dataset are provided in the supplementary materials.
This ability to resolve intra-class variation is a key strength of MAPLE. By encouraging local neighborhoods to align with lower-rank structures through a self-supervised objective, MAPLE introduces perceptible gaps in the final layout between visually distinct sub-groups. From an information visualization perspective, these gaps align with the Gestalt law of proximity [7294427, bae2025uncovering], supporting the perception of distinct groups and facilitating the identification of subtle data subtypes. While we do not claim that all geometric structures in the layout directly correspond to visually identifiable semantic components, the observed consistency between local layout structures and input-level visual patterns suggests that a substantial portion of the internal structures reflects coherent variation in the data and arises from the objective’s suppression of redundant variance. Overall, these results indicate that MAPLE enables higher-resolution visual exploration and supports the discovery of fine-grained intra-class patterns that are often obscured in other DR methods.
@@ IV-C Scalability and Runtime Analysis
We evaluate the performance of MAPLE on a laptop equipped with an Apple M1 Pro chip and 16 GB of unified memory. The training framework is implemented in PyTorch using the Metal Performance Shaders (MPS) backend to execute tensor operations on the integrated GPU. To improve throughput, we use automatic mixed precision for gradient computations. Due to current limitations in the MPS backend for certain singular-value decomposition operators, the nuclear norm required by the MMCR loss is computed on the CPU. To accommodate large datasets under limited memory, we use a batch-loading strategy that avoids memory overflow.
@@
Figure 5: Completion time of MAPLE and UMAP in seconds. Left: runtime with varying feature dimensions ($D$) at fixed $N=10,000$. Right: runtime with varying data sizes (NN) at fixed $D=784$. Solid and dotted lines indicate neighborhood sizes of $k=15$ and $k=30$, respectively. Both axes are shown on logarithmic scales, with tick labels $10^{2}$ and $10^{3}$ indicating orders of magnitude.
Observation.  The results are shown in Fig. 5. MAPLE exhibits approximately linear runtime scaling with respect to both data size $N$ and feature dimensionality $D$, closely matching the scaling behavior of standard UMAP. For instance, increasing the data size from N=50, [^0] N=50{,}000 to N=100, [^0] N=100{,}000 increases runtime by approximately 2.7 times (from roughly 4 minutes to 11 minutes). Runtime also scales linearly with $D$. However, we observe increased sensitivity to neighborhood size for high-dimensional inputs. Beginning at D=2,500D=2{,}500, increasing $k$ from 15 to 30 leads to a significant runtime increase (e.g., from 261 seconds to 1,925 seconds at D=10, [^0] D=10{,}000). This behavior arises from two factors. First, the encoder processes entire local neighborhoods, causing the dominant computational cost to scale as $O(kND^{\prime})$. Second, at high dimensionality, the resulting dense tensors exceed CPU cache capacity, making system memory bandwidth a limiting factor. While batching prevents memory overflow, the runtime bottleneck shifts to data loading and computation. The bottleneck can be partially mitigated by reducing the batch size, although this introduces a trade-off with training efficiency [tsai2021selfsupervised].
The runtime difference between MAPLE and UMAP reflects an inherent trade-off in learning-based DR methods. MAPLE incurs additional computational cost due to the neural network training phase, resulting in a runtime of the order of minutes rather than seconds. However, this cost is predictable and scales linearly with dataset size. In practice, MAPLE’s runtime is comparable to other parametric DR methods [sainburg2021parametric]. We advise practitioners to be mindful of feature dimensionality and neighborhood size when applying MAPLE, particularly in high-dimensional settings. Our results show that MAPLE can process datasets with up to 100,000100{,}000 samples on a standard laptop within minutes, without requiring high-performance hardware.